# Machine-Learning-Enhanced Optimization of Noise-Resilient Variational Quantum Eigensolvers

### v1

@@ 1 Introduction
Particularly relevant for NISQ devices are hybrid quantum-classical algorithms, such as the Variational Quantum Eigensolver (VQE) [5, 6]. VQEs can approximate the ground state of quantum Hamiltonians by classical variational optimization of a parametrized quantum circuit. The quantum computer measures the energy of the quantum system for a given set of parameters, i.e., the objective of the minimization problem, and a classical optimization routine is used to find a better set of parameters. Efficient exploration of the parameter space is crucial for successful optimization.
@@
Despite its efficiency, the NFT algorithm faces several challenges, primarily due to the noise of current quantum hardware. In particular, the algorithm relies on noisy measurements to identify the global minimum in the one-dimensional subspace, which often hinders convergence and complicates optimization. Machine learning offers a promising approach to addressing the challenge of noisy measurements and enhancing the NFT algorithm with more flexibility. In a recent work [9], Nicoli et al. proposed an optimization algorithm called NFT-with-EMICoRe (EMICoRe for short), which extends the NFT framework by incorporating Gaussian Process Regression (GPR) and Bayesian Optimization (BO) for a data-driven and potentially more noise-resilient sequential minimal optimization scheme. However, this work [9] focused solely on measurement shot noise, neglecting the impact of quantum hardware noise, such as decoherence, depolarization, and crosstalk errors.
In these proceedings, we build on prior works [7, 9] and demonstrate that EMICoRe outperforms the NFT baseline, even in the presence of hardware noise. We start in Sec. 2 by reviewing the theoretical foundations of VQE, Gaussian Processes, Bayesian Optimization. In Sec. 3, we introduce the EMICoRe algorithm. This is followed by numerical experiments in Sec. 4, which involve classical simulations of noisy quantum hardware, comparing the results with and without error mitigation. We conclude and provide an outlook in Sec. 5.
@@ 2.1 Variational Quantum Eigensolver (VQE)
The energy $E^{\ast}(\boldsymbol{\theta})$ is the true energy that that would be obtained if the measurement on the quantum computer were error-free. Typically, a quantum computer experiences both Gaussian noise, arising from the statistical nature of repeated measurements (i.e., measurement shot noise), and hardware noise, which results from imperfections in qubits, gates, and measurements [10]. Both types of noise are encoded in the noise term $\varepsilon$ in Eq. 2.
@@
where $\boldsymbol{b}$ is a set of arbitrary coefficients and $\theta_{d}$ is a $D-$dimensional vector of angular parameters for the variational quantum circuit. In particular, this functional form allows the following observation: when keeping all but one parameter constant, i.e.,
@@
the form of the objective function for this one-dimensional subspace of the whole optimization problem reduces to
@@
where $a^{(t)}_{\ell,d}(\ell=1,2,3)$ denote parameters that are independent of $\theta^{(t)}_{d}$, while the superscript $(t)$ refers to the optimization step. As a result, it is enough to measure three points in the one-dimensional space to fit a function over the entire subspace and analytically determine the minimum.
@@
where $\boldsymbol{\hat{\theta}}^{(t-1)}$ denotes the previous best point from step $(t-1)$ [^1] The original paper from Nakanishi et al. [7] proposed a shift $\alpha$ of $\frac{\pi}{2}$ for Eq. 7. In the work by Nicoli et al. [9] it was suggested, that equidistant points for probing the subspace should be used, i.e., a shift of $\frac{2\pi}{3}$ should be preferred. Later, it was derived by Anders et al. [11] that the offset of $\frac{2\pi}{3}$ provides a constant variance over the whole 1D-subspace. In a private communication with the authors, it was reported that, because of a bug, the actual value used in the experiments of [9] was $\frac{\pi}{3}$ instead of $\frac{2\pi}{3}$ as reported in the paper. Although this did not lead to significant differences in the numerical experiments, in order to be consistent with the prior work, we will stick to the convention of using $\frac{\pi}{3}$ for NFT.. This approach is advantageous because it enables analytical optimization of each parameter within its respective one-dimensional subspace, eliminating the need to compute gradients as required by gradient-based methods, such as SPSA [12, 13]. Furthermore, gradient-based approaches may optimize all parameters simultaneously during each gradient update but do not guarantee analytical minimization.
@@ 2.2 Gaussian Processes and Bayesian Optimization
The NFT algorithm can find the exact analytical minimum within the one-dimensional subspace if and only the measurements on the quantum computer are exact, i.e., an infinite number of shots and no hardware noise. However, this is not a real-world scenario. In fact, due to the unavoidable hardware noise on current NISQ hardware, and the inherent statistical uncertainty from repeated measurements (shots), it may be beneficial to enhance existing algorithms with frameworks like Bayesian Optimization (BO) [14] and Gaussian Process Regression (GPR) [15]. These techniques, widely established in machine learning, are often well-suited to deal with noisy data [15]. Let $E^{\ast}(\cdot):\mathcal{X}\mapsto\mathbb{R}$ be an unknown (black-box) objective function to be minimized (i.e., the energy of a VQE in our case). With the aid of a function that suggests promising new points to observe, often referred to as acquisition function [14], one iteratively approximates the target objective with a surrogate function $E(\boldsymbol{\theta})\approx E^{\ast}(\boldsymbol{\theta})$, sampled from the GP. At each optimization step, after new points are measured, those are added to the previous set of measured points, used to update the GP.
A common choice as a surrogate model is a GP regression model with a one-dimensional Gaussian likelihood and GP prior [9]
@@
Here, $\sigma^{2}$ is the variance of the measurement noise $\varepsilon$, and the prior mean function $\nu(\cdot)$ is set to 00 in this work. In the case of VQEs a zero-mean prior is a reasonable choice because the target function $E^{\ast}(\boldsymbol{\theta})$ one aims to model is a linear combination of functions centered around zero (up to a potential constant offset which stems from Eq. 4). The function $k(\cdot,\cdot)$ is the kernel (covariance) function encoding the expected correlation between two measured points. We refer to  Section 3.1 for more details on the kernel function we use in this proceeding. The GP is trained on noisy measurements $\tilde{E}=E^{\ast}(\boldsymbol{\theta})+\varepsilon$ that were collected throughout previous optimization steps.
@@
Let $\{\boldsymbol{\Theta},\boldsymbol{\tilde{E}}\}$ be $N$ training samples, where $\boldsymbol{\Theta}=(\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{N})\in\mathcal{X}^{N}$ are the angular inputs of the parametrized quantum circuit, and $\boldsymbol{\tilde{E}}=(\tilde{E}_{1},\ldots,\tilde{E}_{N})^{\top}\in\mathbb{R}^{N}$ are the energies measured for the corresponding angular parameters $\boldsymbol{\theta}_{i}\in\boldsymbol{\Theta}$. The posterior of a GP regression model, see Eq. 8, is still a GP, i.e., $p(E(\cdot)\mid\boldsymbol{\Theta},\boldsymbol{\tilde{E}})=\operatorname{GP}\left(E(\cdot);\mu_{\boldsymbol{\Theta}}(\cdot),s_{\boldsymbol{\Theta}}(\cdot,\cdot)\right)$. Thus, in Ref. [9] it was shown that for arbitrary $M$ test points $\boldsymbol{\Theta}^{\prime}=\left(\boldsymbol{\theta}_{1}^{\prime},\ldots,\boldsymbol{\theta}^{\prime}{}_{M}\right)\in\mathcal{X}^{M}$, the posterior, or predictive distribution of the function values $\boldsymbol{E}^{\prime}=\left(E\left(\boldsymbol{\theta}_{1}^{\prime}\right),\ldots,E\left(\boldsymbol{\theta}_{M}^{\prime}\right)\right)^{\top}\in\mathbb{R}^{M}$ is the $M$-dimensional Gaussian with its mean and covariance analytically given
@@
$\begin{gathered}p\left(\boldsymbol{E}^{\prime}\mid\boldsymbol{\Theta},\boldsymbol{\tilde{E}}\right)=\mathcal{N}_{M}\left(\boldsymbol{E}^{\prime};\boldsymbol{\mu}_{\boldsymbol{\Theta}}^{\prime},\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}\right),\quad\text{ where }\\ \boldsymbol{\mu}_{\boldsymbol{\Theta}}^{\prime}=\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{\tilde{E}},\quad\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}=\boldsymbol{K}^{\prime\prime}-\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{K}^{\prime}.\end{gathered}$ (9)
@@
The notation with a prime refers to new, unseen, points with the corresponding energies predicted by the GP. Here, $\boldsymbol{K}=k(\boldsymbol{\Theta},\boldsymbol{\Theta})\in\mathbb{R}^{N\times N},\boldsymbol{K}^{\prime}=k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{N\times M}$, and $\boldsymbol{K}^{\prime\prime}=k\left(\boldsymbol{\Theta}^{\prime},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{M\times M}$ are the train, train-test, and test kernel matrices, respectively, where $k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)$ denotes the kernel matrix evaluated at each column of $\boldsymbol{\Theta}$ and $\boldsymbol{\Theta}^{\prime}$ such that $\left(k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\right)_{n,m}=k\left(\boldsymbol{\theta}_{n},\boldsymbol{\theta}_{m}\right)$. Moreover, $\boldsymbol{I}_{N}\in\mathbb{R}^{N\times N}$ denotes the identity matrix. The subscript $\boldsymbol{\Theta}$ in Eq. 9 denotes which input points were used to train the GP. Note that the predictive mean and covariance are given in their vector or matrix form, respectively. The predictive mean can be computed as a function of one arbitrary input vector $\mu_{\boldsymbol{\Theta}}(\boldsymbol{\theta}^{\prime})=\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{\tilde{E}}$ with $\boldsymbol{K}^{\prime}=k\left(\boldsymbol{\Theta},\boldsymbol{\theta}^{\prime}\right)\in\mathbb{R}^{N}$. The computation of a single predictive covariance $s_{\boldsymbol{\Theta}}(\cdot,\cdot)$ follows analogously. For more details on how to compute the predictive mean and variance we refer to Ref. [9].
@@
The knowledge gathered by the GP during optimization, as the number of measurements increases, can not only be exploited for regression but can also be leveraged to choose the next points to observe. This is achieved by combining GPR and BO and solving the following maximization problem
$\max_{\boldsymbol{\Theta^{\prime}}}{a_{\boldsymbol{\Theta}^{(t-1)}}(\boldsymbol{\Theta^{\prime}})}\,$
@@
where $a_{\boldsymbol{\Theta}^{(t-1)}}(\cdot)$ is the so-called acquisition function. This latter uses a GP model trained on previously observed data pairs $\{\boldsymbol{\Theta}^{(t-1)},\boldsymbol{\tilde{E}}\}$, collected over the prior $(t-1)$ iterations, to assess how promising a set of new input points $\boldsymbol{\Theta^{\prime}}$ may be. Points that maximize the acquisition function are likely to be the most informative points to measure at the next iteration, such that the energy is minimized.
@@ 3 Method
In this section, we leverage the functional form of the VQE objective as derived in [7] and we construct a physics-informed kernel suited for GPR. Furthermore, we recall an innovative type of acquisition function and the concept of confident region (CoRe) as introduced in [9]. Finally, in Section 3.2.1, we show how the NFT baseline algorithm can be improved to achieve a more flexible and noise-resilient algorithm when combined with BO and GPR.

### v2

@@ 1 Introduction
Particularly relevant for NISQ devices are hybrid quantum-classical algorithms, such as the Variational Quantum Eigensolver (VQE) [5, 6]. VQEs can approximate the ground state of quantum Hamiltonians by classical variational optimization of a parametrized quantum circuit. The quantum computer measures the energy of the quantum system for a given set of parameters, i.e., the objective of the minimization problem, while a classical optimization routine is used to find a better set of parameters. Efficient exploration of the parameter space is crucial for successful optimization.
@@
Despite its efficiency, the NFT algorithm faces several challenges, primarily due to the noise of current quantum hardware. In particular, the algorithm relies on noisy measurements to identify the global minimum in the one-dimensional subspace, which often hinders convergence and complicates optimization. Machine learning offers a promising approach to addressing the challenge of noisy measurements and enhancing the NFT algorithm with more flexibility. In a recent work [9], Nicoli et al. proposed an optimization algorithm called NFT-with-EMICoRe (EMICoRe for short), which extends the NFT framework by incorporating Gaussian Process Regression (GPR) and Bayesian Optimization (BO) for a data-driven and potentially more noise-resilient SMO scheme. However, this work [9] focused solely on measurement shot noise, neglecting the impact of quantum hardware noise, such as decoherence, depolarization, and crosstalk errors.
In these proceedings, we build on prior works [7, 9] and demonstrate that EMICoRe outperforms the NFT baseline in the presence of simulated quantum hardware noise. We start in Sec. 2 by reviewing the theoretical foundations of VQE, Gaussian Processes, Bayesian Optimization. In Sec. 3, we introduce the EMICoRe algorithm. This is followed by numerical experiments in Sec. 4, which involve classical simulations of noisy quantum hardware, comparing the results with and without error mitigation. We conclude and provide an outlook in Sec. 5.
@@ 2.1 Variational Quantum Eigensolver (VQE)
The energy $E^{\ast}(\boldsymbol{\theta})$ is the true energy that would be obtained if the quantum computer were noise-free. Typically, a quantum computer is subject to both Gaussian noise, arising from the statistical nature of repeated measurements (i.e., measurement shot noise), and hardware noise, which stems from imperfections in qubits, gates, and measurements [10]. Both types of noise are encoded in the noise term $\varepsilon$ in Eq. 2.
@@
where $\boldsymbol{b}$ is a set of arbitrary coefficients and $\theta_{d}$ is a $D$-dimensional vector of angular parameters for the variational quantum circuit. In particular, this functional form leads to the following observation: when keeping all but one parameter constant, i.e.,
@@
the form of the objective function for this one-dimensional subspace of the whole optimization problem reduces to [7]
@@
where $a^{(t)}_{\ell,d}(\ell=1,2,3)$ denote parameters that are independent of $\theta^{(t)}_{d}$, while the superscript $(t)$ refers to the optimization step. As a result, measuring three points in the one-dimensional subspace is sufficient to fit a function across the entire domain and analytically determine its minimum.
@@
where $\boldsymbol{\hat{\theta}}^{(t-1)}$ denotes the previous best point from step $(t-1)$. [^1] The original paper from Nakanishi et al. [7] proposed a shift $\alpha$ of $\frac{\pi}{2}$ for Eq. 7. In the work by Nicoli et al. [9] it was suggested that equidistant points for probing the subspace should be used, i.e., a shift of $\frac{2\pi}{3}$ should be preferred. Later, it was derived by Anders et al. [11] that the offset of $\frac{2\pi}{3}$ provides a constant variance over the whole 1D-subspace. In a private communication with the authors, it was reported that, because of a bug, the actual value used in the experiments of [9] was $\frac{\pi}{3}$ instead of $\frac{2\pi}{3}$ as reported in the paper. Although this did not lead to significant differences in the numerical experiments, in order to be consistent with the prior work, we will stick to the convention of using $\frac{\pi}{3}$ for NFT. This approach is advantageous because it enables analytical optimization of each parameter within its respective one-dimensional subspace, eliminating the need to compute gradients as required by gradient-based methods, such as SPSA [12, 13]. Furthermore, gradient-based approaches may optimize all parameters simultaneously during each gradient update but do not guarantee analytical minimization.
@@ 2.2 Gaussian Processes and Bayesian Optimization
The NFT algorithm can find the exact analytical minimum within the one-dimensional subspace if and only if the state preparation and measurements on the quantum computer are exact—that is, if an infinite number of shots are taken and no hardware noise is present. In a real-world scenario, due to the unavoidable hardware noise on current NISQ hardware and the inherent statistical uncertainty from repeated measurements (shots), it may be beneficial to enhance existing algorithms with frameworks like Bayesian Optimization (BO) [14] and Gaussian Process Regression (GPR) [15]. These techniques, widely established in machine learning, are often well-suited to deal with noisy data [15]. Let $E^{\ast}(\cdot):\mathcal{X}\mapsto\mathbb{R}$ be an unknown (black-box) objective function to be minimized (i.e., the energy in our case). With the aid of a function that suggests promising new points to observe, often referred to as acquisition function [14], one iteratively approximates the target objective with a surrogate function $E(\boldsymbol{\theta})\approx E^{\ast}(\boldsymbol{\theta})$, sampled from the GP. At each optimization step, after new points are measured, those are added to the previous set of measured points, used to update the GP.
A common choice for a surrogate model is a GP regression model with a one-dimensional Gaussian likelihood and a GP prior [9],
@@
Here, $\sigma^{2}$ is the variance of the noise $\varepsilon$ defined in Eq. 2, and the prior mean function $\nu(\cdot)$ is set to 00 in this work. In the case of VQEs, a zero-mean prior is a reasonable choice because the target function $E^{\ast}(\boldsymbol{\theta})$, which one aims to model, is a linear combination of functions centered around zero (up to a potential constant offset stemming from Eq. 4). The function $k(\cdot,\cdot)$ is the kernel (covariance) function encoding the expected correlation between two measured points. We refer to  Section 3.1 for more details on the kernel function we use in these proceedings. The GP is trained on noisy measurements $\tilde{E}=E^{\ast}(\boldsymbol{\theta})+\varepsilon$ that were collected throughout previous optimization steps.
@@
Let $\{\boldsymbol{\Theta},\boldsymbol{\tilde{E}}\}$ be $N$ training samples, where $\boldsymbol{\Theta}=(\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{N})\in\mathcal{X}^{N}$ are the angular inputs of the parametrized quantum circuit, and $\boldsymbol{\tilde{E}}=(\tilde{E}_{1},\ldots,\tilde{E}_{N})^{\top}\in\mathbb{R}^{N}$ are the energies measured for the corresponding angular parameters $\boldsymbol{\theta}_{i}\in\boldsymbol{\Theta}$. The posterior of a GP regression model, see Eq. 8, is still a GP, i.e., $p(E(\cdot)\mid\boldsymbol{\Theta},\boldsymbol{\tilde{E}})=\operatorname{GP}\left(E(\cdot);\mu_{\boldsymbol{\Theta}}(\cdot),s_{\boldsymbol{\Theta}}(\cdot,\cdot)\right)$. Thus, in Ref. [9], it was shown that for arbitrary $M$ test points $\boldsymbol{\Theta}^{\prime}=\left(\boldsymbol{\theta}_{1}^{\prime},\ldots,\boldsymbol{\theta}^{\prime}{}_{M}\right)\in\mathcal{X}^{M}$, the posterior, or predictive distribution of the function values $\boldsymbol{E}^{\prime}=\left(E\left(\boldsymbol{\theta}_{1}^{\prime}\right),\ldots,E\left(\boldsymbol{\theta}_{M}^{\prime}\right)\right)^{\top}\in\mathbb{R}^{M}$, is an $M$-dimensional Gaussian,
@@
$\begin{gathered}p\left(\boldsymbol{E}^{\prime}\mid\boldsymbol{\Theta},\boldsymbol{\tilde{E}}\right)=\mathcal{N}_{M}\left(\boldsymbol{E}^{\prime};\boldsymbol{\mu}_{\boldsymbol{\Theta}}^{\prime},\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}\right),\end{gathered}$ (9)
with mean and covariance analytically given by
$\begin{gathered}\boldsymbol{\mu}_{\boldsymbol{\Theta}}^{\prime}=\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{\tilde{E}},\quad\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}=\boldsymbol{K}^{\prime\prime}-\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{K}^{\prime}.\end{gathered}$ (10)
@@
The notation with a prime refers to new, unseen points with the corresponding energies predicted by the GP. Here, $\boldsymbol{K}=k(\boldsymbol{\Theta},\boldsymbol{\Theta})\in\mathbb{R}^{N\times N},\boldsymbol{K}^{\prime}=k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{N\times M}$, and $\boldsymbol{K}^{\prime\prime}=k\left(\boldsymbol{\Theta}^{\prime},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{M\times M}$ are the train, train-test, and test kernel matrices, respectively, where $k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)$ denotes the kernel matrix evaluated at each column of $\boldsymbol{\Theta}$ and $\boldsymbol{\Theta}^{\prime}$, such that $\left(k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\right)_{n,m}=k\left(\boldsymbol{\theta}_{n},\boldsymbol{\theta^{\prime}}_{m}\right)$. Moreover, $\boldsymbol{I}_{N}\in\mathbb{R}^{N\times N}$ denotes the identity matrix. The subscript $\boldsymbol{\Theta}$ in Eq. 9 denotes which input points were used to train the GP. Note that the predictive mean and covariance are given in their vector and matrix form, respectively. The predictive mean can be computed as a function of one arbitrary test point $\boldsymbol{\theta}^{\prime}$, i.e., $\mu_{\boldsymbol{\Theta}}(\boldsymbol{\theta}^{\prime})=\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{\tilde{E}}$ with $\boldsymbol{K}^{\prime}=k\left(\boldsymbol{\Theta},\boldsymbol{\theta}^{\prime}\right)\in\mathbb{R}^{N}$. The computation of a single predictive covariance $s_{\boldsymbol{\Theta}}(\cdot,\cdot)$ with $\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}=s_{\boldsymbol{\Theta}}\left(\boldsymbol{\Theta}^{\prime},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{M\times M}$ follows analogously. For more details on how to compute the predictive mean and variance, we refer to Ref. [9].
@@
As the number of measurements increases, the knowledge gained by the GP during optimization can be used not only for regression but also to choose the next points to observe. This is achieved by combining GPR and BO and solving the following maximization problem:
$\max_{\boldsymbol{\Theta^{\prime}}}{a_{\boldsymbol{\Theta}^{(t-1)}}(\boldsymbol{\Theta^{\prime}})}\,,$
@@
where $a_{\boldsymbol{\Theta}^{(t-1)}}(\cdot)$ is the so-called acquisition function. This function uses a GP model trained on previously observed data pairs $\{\boldsymbol{\Theta}^{(t-1)},\boldsymbol{\tilde{E}}\}$, collected over the prior $(t-1)$ iterations, to assess how promising a set of new input points $\boldsymbol{\Theta^{\prime}}$ may be. The points that maximize the acquisition function are likely to be the most informative for the next iteration.
@@ 3 Method
In Section 3.1, we leverage the functional form of the VQE objective [7] to construct a physics-informed kernel suited for GPR [9]. In Section 3.2, we introduce a novel type of acquisition function called Expected Maximum Improvement over Confident Regions (EMICoRe) [9]. In Section 3.2.1, we combine these concepts and present the NFT-with-EMICoRe algorithm, called EMICoRe algorithm for short [9]. This algorithm demonstrates how the NFT baseline algorithm can be improved to achieve a more flexible and noise-resilient algorithm when combined with BO and GPR.